Fermi National Laboratory


Accelerator Update for July 8, 2005 - July 11, 2005

Friday July 8
The day shift began with Operations monitoring store 4254 and the Collider Detector at Fermilab (CDF), D-Zero (D0) using the luminosity, with all other beam off due to work on a Linac RF station (LRF4).

A Linac technician reported at 8:34 AM that he had finished his work on LRF4. The Antiproton Source (Pbar) resumed stacking.

Operations established beam to the Meson experiment MCenter at 8:49 AM.

At 10:49 AM LRF4 tripped off. A Linac tech had the station back up within ten minutes.

Operations reported that experts began studies in the MI8 transfer line, which runs between the Booster and the Main Injector (MI), and in the Pbar Accumulator at 3:25 PM. The MI8 studies inhibited all downstream beam.

LRF4 tripped off at 3:59 PM. It was back up by 4:12 PM, but tripped off twice more over the next hour.

Cryo system techs at CHL reported that a Tevatron (TeV) sector B0 compressor had tripped off. The system profile looked good, so nothing was done.

At 6:27 PM, LRF4 tripped for the first of three times in about forty-five minutes. It reset okay for the first two times. It took about twenty-five minutes to bring the current down and the voltage up.

Operations reported that the MI8 studies ended around 7:54 PM.

Operations established beam to NuMI at 8:08 PM.

At 8:15 PM, LRF4 tripped off for the first of three times in forty minutes. It reset okay each time.

Pbar studies ended and stacking resumed at 9:56 PM, but it was poor quality. Operators called in experts who found a loose connection and a bad control card about twenty minutes later.

At 9:57 PM, LRF4 tripped off for the first of two times in about ten minutes. It reset okay.

At 10:52 PM, Pbar cooling went down due to trouble with its TWTs (Traveling Wave Tubes), which made stacking difficult. An Operator called in an expert.

Operations put the Recycler into shot setup at 11:04 to stash antiprotons.

At 11:25 PM, LRF4 tripped off for the first of two times in thirteen minutes. It reset okay.

Saturday July 9
The midnight shift began with Operations monitoring store 4254, with Pbar stacking poorly, with the Recycler in shot setup, and with the following experiments taking beam: CDF, D0, MiniBooNE, and NuMI.

At 12:23 AM, LRF4 tripped off. It reset okay.

At 1:09 AM, a Booster RF station (BRF12) tripped off for the first of four times in an hour and twenty minutes. It reset each time.

At 1:35 AM, LRF4 tripped off for the first of four times in thirty minutes. It reset each time.

Operations reported at 1:47 AM that the Recycler shot setup was on hold due to an Accumulator ARF4 VME problem. An operator called in an expert.

The Pbar expert repaired the ARF4 VME. The Recycler shot was canceled and stacking resumed at 2:42 AM.

MCenter resumed taking beam at 2:54 AM.

At 4:15 AM, Operations reported a problem with a Time Line Generator (TLG) event since Pbar had resumed stacking. Operators continue their investigations. Unfortunately, this problem interfered with MCenter beam.

At 8:31 AM, LRF4 tripped off for the first of four times in two hours. It reset each time. MIRF14 seems to trip off occasionally when LRF4 trips off.

Operations terminated store 4254 at 11:04 PM.

Operations put the TeV into shot setup at 11:15 AM.

Cryo system tech restarted the TeV sector B0 compressor at 11:22 PM.

At 1:48 PM, Operations reported having problems with Pbar's reverse proton efficiency.

Operations established store 4255 at 3:59 PM with an initial luminosity of 69.95E30. Pbar resumed stacking.

MCenter resumed taking beam at 4:17 PM.

At 4:44 PM, LRF4 tripped off for the first of three times in two hours and fifteen minutes. It reset each time.

At 7:30 PM, LRF4 tripped off for the first of three times in thirteen minutes. It reset each time.

At 8:25 PM, LRF4 tripped off. It reset okay.

Operations put the Recycler into shot setup at 8:33 PM to stash antiprotons.

At 8:59 PM, LRF4 tripped off for the first of four times in sixteen minutes. MIRF14 tripped off each time LRF4 tripped off. They both reset each time.

At 9:02 PM, LRF4 tripped off, followed by MRF14, for the first of three times in thirteen minutes. It reset each time.

Stashing ended at 10:15 PM and Pbar resumed stacking.

MCenter resumed taking beam at 10:25 PM.

Sunday July 10
The midnight shift began with Operations monitoring store 4255 and a Recycler stash, with Pbar stacking, and with CDF, D0, MiniBooNE, NuMI, and MCenter taking beam.

At 1:46 AM, LRF4 tripped off. It reset okay.

At 4:22 AM, LRF4 tripped off for the first of four times in about ninety minutes. MIRF14 tripped off each time also. They both reset okay each time.

At 7:59 AM, Operations reported that the Pbar target had stopped rotating. Antiproton production hadnít shown any decrease yet. Operations notified an expert.

At 9:19 AM, LRF4 tripped off. It reset okay.

At 9:29 AM, another Linac RF station (KRF1) tripped off due to LCW cooling problem, a flow valve wasnít responding. This shut off all downstream beam. Operations called in experts who came in and changed out a relay for the system

A Pbar expert reported at 12:13 PM that he needed to access the beam line to fix the antiproton target.

Operations terminated store 4255 at 3:02 PM. Operations put the TeV into shot setup at 3:18 PM.

At 6 PM, the TeV suffered a sector F1 quench due to injecting beam on top of existing beam.

Operations reestablished beam to NuMI at 7:13 PM.

At 8:10 PM, a quadrupole power supply (C:BQ9) tripped off on a quench indication. The TeV was in a dry squeeze at the time.

At 8:23 PM, LRF4 tripped off, followed by MIRF14. They both reset okay.

Operators reported at 9:40 PM that they had replaced the quench detector card for BQ9.

Shot setup resumed at 10:37 PM.

At 11:10 PM, the TeV abort dropped out due to a sector B0 link timeout. Operators replaced a crate power supply.

At 11:24 PM, Operations reported that the main MI LCW valves from CUB closed. Operations informed the Duty Mechanic. The MI system automatically switched to internal recirculation. An expert said that this would be okay until Monday morning.

At 11:32 PM, LRF4, followed by MIRF14, tripped off. They reset okay.

Monday July 11
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

At 12:12 AM, Operations reported that a Lambertson (I:LAM62) tripped off on a ground fault indication. An operator contacted an expert.

At 12:16 AM, LRF4 tripped off. It reset okay.

At 1:23 AM, Operations reported that the Pbar stack was lost due to a ground fault on A:IB. Operations informed everyone.

Operations prepared the Pbar rings for access at 2 AM.

A Pbar power supply expert began work on A:IB at 2:22 AM. At 4 AM he reported that the ground fault was on cables going the to main transformer. Electricians would have to make the repairs.

The Cryo Coordinator said at 7:43 AM that they would use this time to repair a TeV B0 cold compressor, but that the TeV would have to be off for this work. Operators turned off the B0 power supplies.

Operations reported that around 8 AM the electricians had started working on the A:IB cabling.

Around 8 AM, Recycler experts conducted tests that showed the first proof of electron cooling.

The Plan for Monday
The plans for today are to finish the A:IB cabling work, the TeV cryo work, and some miscellaneous TeV maintenance. Hopefully, this will all be completed by noon and, sometime shortly after that, Pbar will resume stacking. TeV and Booster experts will then be allowed to conduct studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/5/2005   email Fermilab